Three times a first number is 2 more than the opposite of a second number. The difference of four times the first number and the second number is 5. Find the smaller of the two numbers.
Answer by mananth(16949)

let the numbers be x & y
3x=2-y
3x+y =2..................1
4x-y=5...................2
Solve for x & y
3 x + y = 2 .............1

4 x - y = 5 .............2
add the two equations

3 x + y = 2
4 x - y = 5
Add the two equations
7x = 7
/ 7
x = 1
plug value of x in (1)
3 x + y = 2
3 + y = 2
y = 2 -3
y = -1
y = -1
-1 is the smaller number